# Logarithmic Regret for Online KL-Regularized Reinforcement Learning

### v3

@@ 5.1 The Proposed Algorithm and Result
6:        Construct bonus $b_{h}^{t}(\cdot,\cdot)$ defined in (5.1).
@@
$\mathcal{F}_{h}^{t-1}=\Big{\{}f_{h}\in\mathcal{F}_{h}:\sum_{i=1}^{t-1}(f_{h}(s_{h}^{i},a_{h}^{i})-\widehat{f}_{h}^{t}(s_{h}^{i},a_{h}^{i}))^{2}+\lambda\leq(\beta_{h}^{T})^{2}\Big{\}},$
@@
$b_{h}^{t}(s,a)=\min\Big{\{}1,\beta_{h}^{T}\cdot U^{h}_{\mathcal{F}_{h}^{t-1}}(\lambda,s,a;\mathcal{D}^{h}_{t-1})\Big{\}},$ (5.1)
@@
$V_{1}^{\pi^{*}}(s_{1})-V_{1}^{\widehat{\pi}}(s_{1})\leq\widehat{V}_{1}(s_{1})-V_{1}^{\widehat{\pi}}(s_{1})$ (5.2)
$\leq\mathbb{E}^{\widehat{\pi}}\big{[}\widehat{Q}_{1}(s_{1},a_{1})-{\mathcal{T}}^{1}\widehat{V}_{2}(s_{1},a_{1})+{\mathcal{T}}^{1}\widehat{V}_{2}(s_{1},a_{1})-{\mathcal{T}}^{1}V_{2}^{\widehat{\pi}}(s_{1},a_{1})\big{]}$
@@
In this work, to avoid the direct summation of Bellman errors as in (5.2), we develop a novel decomposition to obtain the square of the Bellman errors. To realize this goal, instead of decomposing the value functions, we decompose the policies.
@@
where the inequality holds by applying (4.6) with $R^{*}=Q_{h+1}^{\pi^{*}}$ and $\widehat{R}_{t}+b_{t}=\widehat{Q}_{h+1}$. Therefore, we can obtain the following bound for each $(I_{h+1})$.
@@
$I_{h+1}\leq\eta\mathbb{E}^{\widehat{\pi}}_{\cdot|s_{h+1},a_{h+1}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}^{2}\Big{]}.$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $\leq\eta\sum_{h=0}^{H-1}\mathbb{E}^{\widehat{\pi}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}^{2}\Big{]}\leq\eta H^{2}\mathbb{E}^{\widehat{\pi}}\Big{[}\sum_{h=1}^{H}(e_{h}(s_{h},a_{h}))^{2}\Big{]},$
@@ Appendix A Proofs for KL-Regularized Contextual Bandits
$=\frac{1}{\eta}\log Z_{R^{*}}(x)-\frac{1}{\eta}\log Z_{\widehat{R}_{t-1}+b_{t-1}}(x)+\mathbb{E}_{\pi_{t}}\big{[}\widehat{R}_{t-1}(x,a)-R^{*}(x,a)+b_{t-1}(x,a)\big{]},$ (A.1)
@@
$\Delta(x,R):=$ $-\frac{1}{\eta}\log Z_{R}(x)+\sum_{a\in\mathcal{A}}\frac{\pi_{\mathrm{ref}}(a|x)\cdot\exp(\eta R(x,a))}{Z_{R}(x)}\cdot\bigl{(}R(x,a)-R^{*}(x,a)\bigr{)}$ (A.2)
@@
With the definition of $\Delta(x,a)$, the gap in (A.1) can be written as $\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]$. To analyze the behavior of $\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]$, we now study the gradient of $\Delta(x,R)$. We are interested in the gradient of $\Delta(x,R)$ with respect to $R(x,a)$ for each $a\in\mathcal{A}$.
@@
$\frac{\partial\Delta(x,R)}{\partial R(x,a)}$
$=\frac{\partial\big{[}-\frac{1}{\eta}\log Z_{R}(x)+\mathbb{E}_{\pi_{R}^{\eta}}\big{[}R(x,a)-R^{*}(x,a)\big{]}\big{]}}{\partial R(x,a)}$
@@
where the third equality uses (A.4) and the fourth equality uses (A.5) and (A.6), the second equality and the last equality simply use the linearity of $\partial$ and summation.
@@
$\leq\eta\mathbb{E}_{x\sim d_{0}}\bigl{[}\underbrace{\sum_{a\in\mathcal{A}}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}\cdot\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(\lambda)}\bigr{]},$ (A.3)
@@
$\geq 0.$
Therefore, since $f_{\lambda}(x,a)\leq\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)$ for any $(x,a)\in\mathcal{X}\times\mathcal{A}$, we can obtain that
@@
We further have the following bound for RHS of (A.3):
@@
$\frac{\partial Z_{R}(x)}{\partial R(x,a)}=\eta\pi_{\mathrm{ref}}(a|x)\exp\big{(}\eta R(x,a)\big{)},$ (A.4)
$\frac{\partial\pi^{\eta}_{R}(a|x)}{\partial R(x,a)}=\eta\pi_{R}^{\eta}(a|x)-\eta\pi_{R}^{\eta}(a|x)^{2}.$ (A.5)
@@
$\frac{\partial\pi^{\eta}_{R}(a^{\prime}|x)}{\partial R(x,a)}=-\eta\pi_{R}^{\eta}(a|x)\cdot\pi_{R}^{\eta}(a^{\prime}|x).$ (A.6)
@@
$\sum_{i=1}^{t}(\widehat{R}_{t}(x_{i},a_{i})-R^{*}(x_{i},a_{i}))^{2}\leq 8\log(N_{\mathcal{R}}T/\delta)=\frac{1}{2}\beta_{T}^{2}.$ (A.7)
@@
$R^{*}(x,a)-\widehat{R}_{t}(x,a)$
$\leq\sup_{R_{1},R_{2}\in\mathcal{R}}\frac{|R_{1}(x,a)-R_{2}(x,a)|}{\sqrt{\lambda+\sum_{i=1}^{t}(R_{1}(x_{i},a_{i})-R_{2}(x_{i},a_{i}))^{2}}}\cdot\sqrt{\lambda+\sum_{i=1}^{t}(R^{*}(x_{i},a_{i})-\widehat{R}_{t}(x_{i},a_{i}))^{2}}$
@@
where the second inequality follows from the definition of $U_{\mathcal{R}}$ and (A.7), and the last inequality uses $\lambda\leq\frac{1}{2}\beta_{T}^{2}$ and the choice of bonus $b_{t}$. Besides, since $R^{*}(x,a)-\widehat{R}_{t}(x,a)\leq 1$, we have
@@
where the second inequality uses the condition $\mathcal{E}_{t-1}$, the second equality uses the definition of bonus $b_{t}$ in (4.1), and the last inequality follows from the value of $\beta_{T}$ and the eluder dimension in Definition 3.3. ∎
@@ B.1 Regret Decomposition
$\leq\eta\mathbb{E}^{\widehat{\pi}}_{\cdot|s_{h+1},a_{h+1}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}^{2}\Big{]},$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $\leq\eta\sum_{h=0}^{H-1}\mathbb{E}^{\widehat{\pi}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}^{2}\Big{]}.$
@@
$\sum_{i=1}^{t-1}\big{[}\widehat{f}^{t}_{h}(s_{h}^{i},a_{h}^{i})-{\mathcal{T}}_{\eta}^{h}\widehat{V}^{t}_{h+1}(s_{h}^{i},a_{h}^{i})\big{]}^{2}\leq 8\log(4N_{h}tH/\delta)\leq\frac{1}{2}(\beta^{T}_{h})^{2},$ (B.1)
@@
$\leq(\beta_{h}^{T})^{2}\cdot U^{h}_{\mathcal{F}_{h}}(s,a;\mathcal{D}^{h}_{t-1}),$ (B.2)
where the first inequality holds by using the completeness of the Bellman operator ${\mathcal{T}}_{\eta}^{h}\widehat{V}^{t}_{h+1}\in\mathcal{F}_{h}$ and the definition of uncertainty $U^{h}_{\mathcal{F}_{h}}$, and the last inequality applies (B.1) and $\lambda\leq\frac{1}{2}(\beta^{T}_{h})^{2}$.
@@
$\widehat{Q}^{t}_{h}(s_{h},a_{h})-Q^{*}_{h}(s_{h},a_{h})$
$=\widehat{f}^{t}_{h}(s_{h},a_{h})+b_{h}^{t}(s_{h},a_{h})-{\mathcal{T}}_{\eta}^{h}V_{h+1}^{*}(s_{h},a_{h})$
@@
where the first inequality uses (B.1) and the definition of bonus $b_{h}^{t}$ in (4.1), the second inequality follows from the derivation of the value function in (3.1), and the last inequality is due to the induction hypothesis at step $h+1$. ∎
@@ Appendix D Auxliary Lemmas
For a detailed proof, see Theorem 13.2 of Zhang (2023) or Lemma 34 of Foster and Rakhlin (2023). ∎

### v5

@@ 5.1 The Proposed Algorithm and Result
6:        Construct bonus $b_{h}^{t}(\cdot,\cdot)$ defined in (9).
@@
$\mathcal{F}_{h}^{t-1}=\Big{\{}f_{h}\in\mathcal{F}_{h}:$
$\qquad\qquad\sum_{i=1}^{t-1}(f_{h}(s_{h}^{i},a_{h}^{i})-\widehat{f}_{h}^{t}(s_{h}^{i},a_{h}^{i}))^{2}+\lambda\leq(\beta_{h}^{T})^{2}\Big{\}},$
@@
$b_{h}^{t}(s,a)=\min\Big{\{}1,\beta_{h}^{T}\cdot U^{h}_{\mathcal{F}_{h}^{t-1}}(\lambda,s,a;\mathcal{D}^{h}_{t-1})\Big{\}},$ (9)
@@
$V_{1}^{\pi^{*}}(s_{1})-V_{1}^{\widehat{\pi}}(s_{1})\leq\widehat{V}_{1}(s_{1})-V_{1}^{\widehat{\pi}}(s_{1})$ (10)
$\leq\mathbb{E}^{\widehat{\pi}}\big{[}\widehat{Q}_{1}(s_{1},a_{1})-{\mathcal{T}}^{1}\widehat{V}_{2}(s_{1},a_{1})$
$\qquad+{\mathcal{T}}^{1}\widehat{V}_{2}(s_{1},a_{1})-{\mathcal{T}}^{1}V_{2}^{\widehat{\pi}}(s_{1},a_{1})\big{]}$
@@
In this work, to avoid the direct summation of Bellman errors as in (10), we develop a novel decomposition to obtain the square of the Bellman errors. To realize this goal, instead of decomposing the value functions, we decompose the policies.
@@
where the inequality holds by applying (8) with $R^{*}=Q_{h+1}^{\pi^{*}}$ and $\widehat{R}_{t}+b_{t}=\widehat{Q}_{h+1}$. Therefore, we can obtain the following bound for each $(I_{h+1})$.
@@
$I_{h+1}\leq\eta\mathbb{E}^{\widehat{\pi}}_{\cdot|s_{h+1},a_{h+1}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}\Big{)}^{2}\Big{]}.$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $\leq\eta\sum_{h=0}^{H-1}\mathbb{E}^{\widehat{\pi}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}\Big{)}^{2}\Big{]}$
$\leq\eta H^{2}\mathbb{E}^{\widehat{\pi}}\Big{[}\sum_{h=1}^{H}(e_{h}(s_{h},a_{h}))^{2}\Big{]},$
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
We thank the anonymous reviewers for their helpful comments. HZ and QG are supported in part by the National Science Foundation DMS-2323113 and IIS-2403400. HZ is also partially supported by the Amazon PhD fellowship. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Proofs for KL-Regularized Contextual Bandits
$=\frac{1}{\eta}\log Z_{R^{*}}(x)-\frac{1}{\eta}\log Z_{\widehat{R}_{t-1}+b_{t-1}}(x)+\mathbb{E}_{\pi_{t}}\big{[}\widehat{R}_{t-1}(x,a)-R^{*}(x,a)+b_{t-1}(x,a)\big{]},$ (11)
@@
$\Delta(x,R):=$ $-\frac{1}{\eta}\log Z_{R}(x)+\sum_{a\in\mathcal{A}}\frac{\pi_{\mathrm{ref}}(a|x)\cdot\exp(\eta R(x,a))}{Z_{R}(x)}\cdot\bigl{(}R(x,a)-R^{*}(x,a)\bigr{)}$ (12)
@@
With the definition of $\Delta(x,a)$, the gap in (11) can be written as $\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]$. To analyze the behavior of $\mathbb{E}_{x\sim d_{0}}[\Delta(x,\widehat{R}_{t-1}+b_{t-1})-\Delta(x,R^{*})]$, we now study the gradient of $\Delta(x,R)$. We are interested in the gradient of $\Delta(x,R)$ with respect to $R(x,a)$ for each $a\in\mathcal{A}$.
@@
$\frac{\partial\Delta(x,R)}{\partial R(x,a)}$ $=\frac{\partial\big{[}-\frac{1}{\eta}\log Z_{R}(x)+\mathbb{E}_{\pi_{R}^{\eta}}\big{[}R(x,a)-R^{*}(x,a)\big{]}\big{]}}{\partial R(x,a)}$
@@
where the third equality uses (14) and the fourth equality uses (15) and (16), the second equality and the last equality simply use the linearity of $\partial$ and summation.
@@
$\leq\eta\mathbb{E}_{x\sim d_{0}}\bigl{[}\sum_{a\in\mathcal{A}}\bigl{(}\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)-R^{*}(x,a)\bigr{)}\cdot\underbrace{\bigl{(}f_{\lambda}(x,a)-R^{*}(x,a)\bigr{)}\cdot\pi_{f_{\lambda}}^{\eta}(a|x)}_{U(x,a)}\bigr{]},$ (13)
@@
$\geq 0,$
, which follows from Lemma 2.14 in (Zhao et al., 2025). Therefore, since $f_{\lambda}(x,a)\leq\widehat{R}_{t-1}(x,a)+b_{t-1}(x,a)$ for any $(x,a)\in\mathcal{X}\times\mathcal{A}$, we can obtain that
@@
We further have the following bound for RHS of (13):
@@
$\frac{\partial Z_{R}(x)}{\partial R(x,a)}=\eta\pi_{\mathrm{ref}}(a|x)\exp\big{(}\eta R(x,a)\big{)},$ (14)
$\frac{\partial\pi^{\eta}_{R}(a|x)}{\partial R(x,a)}=\eta\pi_{R}^{\eta}(a|x)-\eta\pi_{R}^{\eta}(a|x)^{2}.$ (15)
@@
$\frac{\partial\pi^{\eta}_{R}(a^{\prime}|x)}{\partial R(x,a)}=-\eta\pi_{R}^{\eta}(a|x)\cdot\pi_{R}^{\eta}(a^{\prime}|x).$ (16)
@@
$\sum_{i=1}^{t}(\widehat{R}_{t}(x_{i},a_{i})-R^{*}(x_{i},a_{i}))^{2}\leq 8\log(N_{\mathcal{R}}T/\delta)=\frac{1}{2}\beta_{T}^{2}.$ (17)
@@
$R^{*}(x,a)-\widehat{R}_{t}(x,a)$ $\leq\sup_{R_{1},R_{2}\in\mathcal{R}}\frac{|R_{1}(x,a)-R_{2}(x,a)|}{\sqrt{\lambda+\sum_{i=1}^{t}(R_{1}(x_{i},a_{i})-R_{2}(x_{i},a_{i}))^{2}}}\cdot\sqrt{\lambda+\sum_{i=1}^{t}(R^{*}(x_{i},a_{i})-\widehat{R}_{t}(x_{i},a_{i}))^{2}}$
@@
where the second inequality follows from the definition of $U_{\mathcal{R}}$ and (17), and the last inequality uses $\lambda\leq\frac{1}{2}\beta_{T}^{2}$ and the choice of bonus $b_{t}$. Besides, since $R^{*}(x,a)-\widehat{R}_{t}(x,a)\leq 1$, we have
@@
where the second inequality uses the condition $\mathcal{E}_{t-1}$, the second equality uses the definition of bonus $b_{t}$ in (3), and the last inequality follows from the value of $\beta_{T}$ and the eluder dimension in Definition 3.3. ∎
@@ B.1 Regret Decomposition
$\leq\eta\mathbb{E}^{\widehat{\pi}}_{\cdot|s_{h+1},a_{h+1}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}\Big{)}^{2}\Big{]},$
@@
$J(\pi^{*})-J(\widehat{\pi})$ $\leq\eta\sum_{h=0}^{H-1}\mathbb{E}^{\widehat{\pi}}\Big{[}\Big{(}\sum_{j=h+1}^{H}e_{j}(s_{j},a_{j})\Big{)}\Big{)}^{2}\Big{]}.$
@@
$\sum_{i=1}^{t-1}\big{[}\widehat{f}^{t}_{h}(s_{h}^{i},a_{h}^{i})-{\mathcal{T}}_{\eta}^{h}\widehat{V}^{t}_{h+1}(s_{h}^{i},a_{h}^{i})\big{]}^{2}\leq 8\log(4N_{h}tH/\delta)\leq\frac{1}{2}(\beta^{T}_{h})^{2},$ (18)
@@
$\leq(\beta_{h}^{T})^{2}\cdot U^{h}_{\mathcal{F}_{h}}(s,a;\mathcal{D}^{h}_{t-1}),$ (19)
where the first inequality holds by using the completeness of the Bellman operator ${\mathcal{T}}_{\eta}^{h}\widehat{V}^{t}_{h+1}\in\mathcal{F}_{h}$ and the definition of uncertainty $U^{h}_{\mathcal{F}_{h}}$, and the last inequality applies (18) and $\lambda\leq\frac{1}{2}(\beta^{T}_{h})^{2}$.
@@
$\widehat{Q}^{t}_{h}(s_{h},a_{h})-Q^{*}_{h}(s_{h},a_{h})$ $=\widehat{f}^{t}_{h}(s_{h},a_{h})+b_{h}^{t}(s_{h},a_{h})-{\mathcal{T}}_{\eta}^{h}V_{h+1}^{*}(s_{h},a_{h})$
@@
where the first inequality uses (B.1) and the definition of bonus $b_{h}^{t}$ in (3), the second inequality follows from the derivation of the value function in (1), and the last inequality is due to the induction hypothesis at step $h+1$. ∎
@@ Appendix D Auxliary Lemmas
For a detailed proof, see Theorem 13.2 of Zhang (2023) or Lemma 34 of Foster & Rakhlin (2023). ∎